OpenStudy (anonymous):

give the minimum value of x if 2x+y is greater then or equal to 7x-9 and y=-3

OpenStudy (anonymous):

2x-3>=7x-9 6>=5x 6/5>=x

OpenStudy (dumbcow):

there is no min value for x the max value is 6/5